### v1

# From the light curves of Long Period Variables to their evolution along the AGB

Abstract
A sample of 116 Long Period Variables is studied with the aim to reveal relations between the properties of their light curves and of their evolution along the Asymptotic Giant Branch (AGB). Each light curve is carefully scrutinized and its properties are summarized in a set of five parameters. The picture that had been sketched in an earlier work is clarified and new features are revealed. In particular, the evolution on the AGB of stars having not yet experienced a third dredge-up event is shown to go together with a closely related path followed by the light curve in the parameter space. Moreover, evidence is found for the existence of two different types of light curves for stars having experienced dredge-up events strong enough for their impact to be detectable. Interpretations are proposed, however often too speculative to be firmly ascertained: the results presented raise several unanswered questions and point to the need for new observations and analyses, suggesting that light curves still carry more messages and information than has been currently possible to decrypt.
show]pttnhung@vnsc.org.vn
show]mainhattanhdpt@gmail.com
show]dthoai@vnsc.org.vn
show]darriulat@vnsc.org.vn
1 Introduction
In a recent article (D. T. Hoai et al., 2025) (Paper I), we presented a detailed study of a sample of light curves of Mira variables selected for a high density of observations, from which we inferred a number of relations between their properties and the evolution of the star along the Asymptotic Giant Branch (AGB). In its wake, the present article clarifies the picture presented by Paper I and addresses questions that were left unanswered by using a larger sample of curves and introducing new parameters.
We refer the reader to Paper I for an introduction to the physics of variable stars in general and of Mira variables in particular. We simply recall that the basic mechanism causing stars to pulsate is the action of a valve allowing or preventing heat to escape, depending on the state of the star. In the case of Cepheids, which were studied first, the valve is essentially the result of the dependence of the opacity of an ionized helium layer on temperature and density (kappa mechanism). In the case of AGB stars (F. Herwig, 2005), the energy transport is dominantly convective rather than radiative, which requires a dynamic theory of convection. Recently, significant progress toward producing realistic three-dimensional models of convection has been achieved (F. Addari et al. 2024 and references therein; B. Freytag & S. Höfner 2023 and references therein) and the role played by convective cells in the evolution of AGB stars has become an important topic of study, both for observations and for modelling. However, as remarked by J. R. Buchler (1993) in the case of Cepheids, numerical simulations, the brute-force method that produces the most detailed and accurate description of observations, fail to reveal simply the underlying mechanism at stake. This is still true today and the same can be said, even more strongly, of Long Period Variables.
Fifty years ago, M. Schwarzschild (1975) argued that convection was probably a major contributor to the mass loss rate of Red Giants and was likely to materialize in the form of a very few convective cells, each covering typically a steradian. His argument was based on a scaling of the Sun’s supergranulation. Today’s understanding of solar supergranulation has somewhat progressed (F. Rincon & M. Rieutord, 2018), but the argument still applies in broad terms and observations using VLTI and ALMA, together with VLBI detections of masers, have confirmed the existence of such convective cells. In broad terms, it seems well established that the mass loss of oxygen-rich stars has its source in radial movements associated with pulsations and the shock waves which they generate, both at the global star and convective cell levels, whether at steradian scale with life times of years or at smaller scales with much shorter life times (C. Paladini et al., 2018; A. Rosales-Guzmán et al., 2024; W. Vlemmings et al., 2017; W. H. T. Vlemmings et al., 2019; W. Vlemmings et al., 2024; P. Darriulat et al., 2024).
But there is more to it. The study of stellar pulsations offers a number of puzzling features that contribute to the frustration of often failing to produce a simple description of what is observed. Much light was shed on these issues during the last quarter of the past century. By then E. N. Lorenz (1964) had studied the logistic map, a recurrence relation of the form and understood how simple non-linear dynamical equations can produce chaos; R. M. May (1976) had realized the strong impact that such results can have on many branches of physical sciences and several authors, such as Alexander Sharkovsky and Mitchell Feigenbaum had studied periodic non-linear dynamical systems (M. J. Feigenbaum, 1978). In astrophysics, during the fourth quarter of the past century, such results were applied to stellar pulsations, mostly Cepheids (J. R. Buchler, 1993). Starting with writing the complicated set of perturbative partial differential equations describing the sources of non-linearity governing the physics of stellar pulsations, one can reduce it to a small number (called number of principal modes) of ordinary differential equations. In the same way as mathematicians had shown how simple mathematics can produce very complicated dynamics, astrophysicists have then shown how complex systems can be described in simple terms. Examples are a non-resonant RR Lyrae model, quite accurately described by the system of two non-resonant amplitude equations for the fundamental mode and the first overtone (J. R. Buchler & G. Kovacs, 1987); the 2:1 resonance between the fundamental mode and the second overtone found to be responsible for the Hertzsprung Progression in bump Cepheids (E. Hertzsprung, 1926; G. Kovacs & J. R. Buchler, 1989); or the 3:2 resonance found to produce alternating cycles and the 5:2 resonance being responsible for period doubling. One may also mention two RV Tauri stars, with light curves displaying double-peaked oscillations and apparent chaos: AC Her (Z. Kollath et al., 1998) and R Sct (J. R. Buchler et al., 1995).
Indeed, many peculiarities observed in the light curves of Long Period Variables are far from being their exclusive and privileged property. Period-doubling is a common phenomenon in many branches of physics, humps and alternately deep and shallow minima are often found in other types of stars. Examples are the above-mentioned RV Tauri stars, W Virginis stars (E. Plachy et al., 2017), BL Herculis stars (R. Smolec et al., 2012; I. Soszyński et al., 2011), RR Lyrae stars (R. Smolec et al., 2015), bump Cepheids (G. Kovacs & J. R. Buchler, 1989), etc. The dynamics of these stars are governed by different mechanisms, hydrogen or helium burning, radiative or convective heat transfer, etc. and such peculiarities are simply general and well-known features of non-linear dynamical systems: their interpretation in terms of specific underlying physical processes is not straightforward. When studying stellar pulsations, one must keep in mind that these are complex dynamical systems, small changes in the parameters governing their dynamics may have major impacts: the light curves are prone to experience strong and/or sudden changes in period, amplitude and regularity. In particular, when seeking interpretations of the results of Fourier or wavelet analyses of light curves displaying complex behaviour, the remarkably simple picture often offered by an analysis of the power spectrum should not be interpreted as evidence for pulsating stars being organ pipes. This has been remarkably illustrated by (A. B. Fokin, 1994), who has produced a most insightful study of the pulsations of AC Her, an RV Tauri star, showing how a sequence of complex events at stake inside the star could be simply described as a resonance between fundamental and first overtone.
2 The sample of light curves and their parameterization
In Paper I, we used two partly overlapping samples of curves, both selected from the list of Mira variables produced by P. Merchán Benítez et al. (2023): Sample A included 71 stars of spectral types C, S and M, the latter restricted to stars of which the spectrum had been searched for technetium with a clear result, positive (referred to as Myes types) or negative (referred to as Mno types). Sample B included 24 curves from sample A selected for a very high density of observations, allowing for a particularly reliable parameterization; in addition, it included 8 curves of similar high quality, of stars of spectral type M of which the technetium content is unknown.
The present sample includes 37 additional stars from the P. Merchán Benítez et al. (2023) selection, of which 7 chosen for having a period in excess of 470 days, and two stars excluded from it as being classified as semi-regular, W Hya and RS Cyg, which both display interesting features. W Hya was studied by S. Uttenthaler et al. (2011) as meander-type period-changing and was found to display no technetium but some lithium in the spectrum, although unlikely to imply efficient Hot Bottom Burning (HBB) operation. RS Cyg is a carbon-rich variable with a period of some 420 days and a low mass loss rate of 210-7 M⊙yr-1 (R. Guandalini et al., 2006), possibly presenting a double-peaking profile (R. R. Cadmus, 2022, 2024).
It is important to recall that, when evolving along the AGB, some stars pulsate, others don’t. Among stars that pulsate, some do it in a very regular way, with well-defined periods and oscillation amplitudes, others display instead significant irregularity, with changing periods and amplitudes. The traditional separation between Mira variables and semi-regular variables is an attempt to distinguish between them, however often failing to do so in a convincing way. In the present article, it is sufficient to be conscious that during their AGB life, most stars spend some time displaying pulsations of sufficient regularity to allow for the definition of a mean period and amplitude; we shall refer to these as being in the domain of stable pulsation; others do not, but there is continuity between the two cases and several stars will be seen to be in a state of transition between the two domains.
We study light curves produced by the American Association of Variable Stars Observers (AAVSO), limited to the past century, from January 1, 1925 to December 31, 2024, and to the visible band. This choice is dictated by the need to deal with curves offering a large enough density of observations to allow for a reliable definition of their shapes. We inspected carefully and separately each of the light curves associated with the sample of stars selected in the present study. Differences from cycle to cycle are often stronger than between mean cycles of different stars, a well-known source of complexity that makes the study of such curves particularly difficult. In particular, many different parameterizations may be devised in order to attempt as reliable and useful a characterization of their shapes. In the present article, we use five such parameters, four of which were defined in Paper I: the mean period , the mean oscillation amplitude , the mean asymmetry between the ascending and descending branches and a parameter measuring the irregularity of the curve. The fifth parameter, , measures the width of the profile in each cycle and is defined below. In addition to these five parameters chosen to characterize the light curve, we study their relation to two parameters that characterize the evolution of the star along the AGB: a colour index C and the spectral type, Mi-j, S or C, both of which are listed in the P. Merchán Benítez et al. (2023) list. We briefly comment on each of these below.
The period is measured in days. Small differences between the values listed by P. Merchán Benítez et al. (2023) and those used in the present article or in Paper I are the result of their being evaluated on different intervals of time and are irrelevant to the present study. Stronger period changes have been observed and studied in the published literature, A. A. Zijlstra & T. R. Bedding (2002) being first to make a distinction between different types of period changes, which has then been generally adopted in later studies (P. Merchán Benítez et al., 2023, and references therein). R Cen and T UMi are commonly quoted as cases of extreme and sudden changes of the periods and oscillation amplitudes of their light curves; we comment on these in Subsections 4.1 and 4.2, respectively. Other examples of such changes, however less spectacular, are LX Cyg, BH Cru and RU Vul. LX Cyg and BH Cru have been shown by S. Uttenthaler et al. (2011, 2016) to have recently experienced a third-dredge-up event and to have switched on this occasion from spectral type S to C; they are part of the present sample. RU Vul is a semi-regular variable of spectral type M4, which has seen its period decrease from 155 to 108 days in the past 55 years (M. Templeton et al., 2008).
Examples of major, but less sudden period changes are R Hya, R Aql and W Dra; they are part of the present sample. R Hya saw its period decrease from 500 days to 360 days over 250 years. P. R. Wood & D. M. Zarro (1981) suggested that such decline was the result of a recent thermal pulse, but A. A. Zijlstra et al. (2002) raised questions in relation with its mass loss history and S. Uttenthaler et al. (2011) considered the possibility of other scenarios. Recently, M. Joyce et al. (2024) and Y. A. Fadeyev (2024) have produced state-of-the-art models of the pulsations; Fadeyev concludes that the best fit is obtained for an initial mass of 4.7 solar masses, and Joyce et al. obtain instead an initial mass of 1.6 to 3.7 solar masses, illustrating the difficulty to produce reliable modelling.
In Paper I we introduced two parameters meant to measure the irregularity displayed by a light curve: and . is the rms deviation from its mean of the magnitude at maximum light, is the mean value of the difference of magnitude between two successive light maxima. We studied the relation between these two parameters and gave evidence for different types of irregularity. However, for the purpose of the arguments presented in the present study, it will be sufficient to use a single parameter, the mean value of and .
A particularly useful parameter largely used in Paper I is the asymmetry between the durations of the ascending and descending branches. In the case of sample B we preferred to use the equivalent parameter . Depending on the method used to measure such asymmetry, different systematic uncertainties are attached to and . In the present article we systematically use the former. We recall that in Paper I, we noted important differences between the properties of the light curves of spectral type Mno according to the value of : we defined a first Mno family, Mno1, as having . In contrast with curves of the second family, Mno2, their light curves were found to display no hump on the ascending branch, to be more symmetric with between 0.50 and 0.55 (compared with between 0.57 and 0.65 for Mno2), to have smaller periods, below 300 days, to be more regular and to have smaller colour indices than Mno2. They sit on the left-hand side of the colour vs period diagrams first produced by S. Uttenthaler et al. (2019), while Mno2 occupies also the right-hand side. When observed in infrared, the oscillation amplitude is less than 20% of the value in the visible, while it exceeds 25% for Mno2.
The mean oscillation amplitude displayed by the light curves, , measured in units of magnitude, was evaluated for each light curve of the Paper I sample, but had been made little use of. In the present study, we pay more attention to the information that it carries.
A well-known feature displayed by the light curves is a clear difference between profiles having a narrow maximum and a broad minimum and profiles having a broad maximum and a narrow minimum. The left and central panels of Figure 1 display two extreme cases. Such distinction had already been made by T. Lebzelter (2011), however without producing very significant results. We searched for parameters that quantify reliably this feature and found two good candidates, which we call and (Figure 1, right); is the width of the curve, relative to the period, at an apparent luminosity equal to 20% of the luminosity amplitude of the oscillation; is the ratio between the luminosity measured from minimum light and integrated over a period and the product of its value at maximum light by the period. Pulsations with relatively short values of the parameters mean a star that is mostly dim during its cycle, pulsations with relatively large values mean a star that is mostly bright during its cycle. We find that, in general, the two parameters are closely related, with, on average, 1.5 . Apart from a small difference between oxygen-rich and carbon-rich stars, we did not find significant cases of deviation between the information provided by the two parameters and considered it sufficient, for the purpose of the present study, to use a single width parameter , defined as the geometric mean between and , . The main difference between parameter and similar parameters commonly used in the published literature, including Paper I, is its being evaluated from the luminosity rather than from the magnitude. This has a trivial but strong effect when the oscillation amplitude is large; a good illustration is the case of chi Cyg which has a very low value of , 0.18, while its profile, studied in Paper I, is broad and displays a hump on the ascending branch.
Finally, we use the colour indices K-[22] and [3.4]-[22] listed by P. Merchán Benítez et al. (2023) to reveal the emission of dust in the circumstellar envelope; the K band is sensitive to water vapour emission (G. Perrin et al., 2015) and the WISE [3.4] band may contain strong absorption features of HCN and/or C2H2 molecules (W. Aoki et al., 1999). P. Merchán Benítez et al. (2023) remarked that the advantage of the [3.4] band is that it was measured at the same time as the [22] band but its disadvantage is that bright Mira variables saturate some detector pixels. We studied the relation obeyed by these parameters and concluded that, for the purpose of the present study, it is sufficient to use a single parameter, (K+[3.4])-[22].
There is a broad spectrum of density of observations among the curves studied in the present article. As a result, the uncertainties attached to the curve parameters cover a broad range. We dealt carefully with each case separately, implying, in several cases, a selection of cycles for which we trusted that a reliable evaluation could be obtained. In particular, several curves are the result of observations made from a limited range of latitudes; when the star is low in the sky, this implies that it can be reliably observed during only part of the year; when the pulsation period is close to a year, the result is that for a long sequence of years a same part of a cycle is observed, producing so-called seasonal gaps: for example, the curve may cover the ascending branch for a decade and the descending branch for the next. We have good control over such difficulties and trust that the uncertainties attached to the curve parameters are small enough to have no significant impact on the arguments and conclusions presented in the article. However, we refrained from quoting uncertainties on the values of each of the measured parameters.
3 Distributions of the light curves in the parameter space
3.1 The Mno spectral type
As was discussed in Paper I, the Mno spectral type is expected to include stars that have recently entered the TP-AGB or, possibly, that are still in the E-AGB and will soon do so. The latter are still burning helium around the degenerate C/O core, while the former are usually burning hydrogen around the depleted helium shell; however, when this has produced enough helium ashes in the inter-shell layer, they experience a brief thermal pulse, which, when strong enough, causes a third dredge-up (TDU) event. Low initial mass stars never experience such TDU event and spend their entire TP-AGB life as Mno types; only higher initial mass stars eventually switch from Mno to Myes spectral type.
In the present section we study the light curves of stars having an Mno spectral type, as listed in Table 1. More precisely, we exclude four such stars that display peculiarities that will be discussed in Subsection 4.2 (W Dra, W Hya, T Ari and U CMi), and we include a star of spectral type M2, RT Cyg, for which technetium had been searched for and found unlikely to be present. Figure 2 illustrates the distributions of the curves in the parameter space by displaying projections on the vs , vs , vs , vs and vs planes. It reveals a remarkable dependence on the temperature of the star, as measured by the value of index for an M- spectral type. As increases, namely as the star cools down, decreases and all other parameters increase. This evolution, for which Table 2 gives evidence, is particularly clear in the cases of , , and , slightly less in the case of and unclear in the case of .
This result confirms and sheds light on the evolution from a first to a second family of curves that was suggested in Paper I. However, it gives evidence for a continuous evolution rather than for a separation between two distinct families. We note that none of the curves listed in Table 1 displays a strong hump on the ascending branch. However, a few late types, such as R Cas, U Her, R LMi and R Leo, do occasionally display a weak hump on the lower part of the ascending branch, in confirmation of the assignment of a b profile type made in Paper I.
This implies that as stars of the Mno spectral type evolve along the AGB, the width parameter of their light curves, , decreases. This raises two questions: where do they come from? And where will they go to?
As strongly suggested by the vs distribution shown in Figure 2, the progenitors of our Mno sample are variables of smaller period, which are absent from the P. Merchán Benítez et al. (2023) selection because Mira variables are required by definition to have periods in excess of 100 days. An interesting contribution to this issue is offered by a detailed study of the circumstellar envelope of L2 Pup by P. Kervella et al. (2014), showing a strong similarity with R Vir, one of the early Mno stars in our sample, although their spectral types are different: M3.5 for R Vir and M5 for L2 Pup. Another very detailed study by S. Uttenthaler (2024) has given evidence for L2 Pup to be Tc-poor and in an early stage of its evolution on the AGB, or even still on the RGB, approaching its tip. Numerous observations at various wavelengths have been reported (for a recent analysis see D. T. Hoai et al., 2022a, and references therein), quoting a low initial mass, 1 to 2 solar masses, and a very low mass loss rate, below M⊙yr-1. L2 Pup seems therefore to be a good candidate for being in a state preceding that of the early Mno types in our sample, such as R Vir or R Vul. Its light curve displays episodes during which the luminosity varies by less than one unit of magnitude, separated by episodes during which it pulsates regularly (0.2), with a mean oscillation amplitude of 1.4 units of magnitude, a stable period of 137 days and a width parameter 0.5. Its being in an unstable regime of pulsation at the limit of what was defined in the introduction as the domain of stable pulsation, suggests that in their early life on the AGB, stars may see their light curves progressively evolve from irregular to Mira-like once having reached the TP-AGB.
When the star has a large enough initial mass, it will at some point experience sufficiently strong TDU for its spectral type to change to Myes and eventually S. If the above interpretation of the evolution of its light curve is correct, we should therefore expect the associated light curves to populate the low part of the parameter space, say below 0.3. In the next subsection, we use this argument as guideline to seek an answer to the second question that we need to address: what will the stars in our Mno sample evolve to?
3.2 Oxygen-rich stars having possibly experienced strong TDU for which 0.3
The study presented in the preceding subsection suggested that the stars of Mno spectral type listed in Table 1, when initially sufficiently massive, would evolve to the Myes spectral type with light curves of small width parameter, say 0.3. Table 3 lists all curves of our sample, excluding those listed in Table 1, that have an M or S spectral type and for which the light curve has a width parameter 0.3. Their distribution in the parameter space is displayed in Figure 3. As decreases, the period, the oscillation amplitude and the colour index tend to increase. Remarkably, a dozen of curves display large irregularity parameters, well in excess of 0.6, in strong contrast with the Mno sample where a single curve, that of R Cas, has a large value of ; it has also extreme values of , and , suggesting that it is close to evolve to an Myes spectral type. The presence of a large group of irregular curves in the present sample suggests that some are evolving toward a state of unstable pulsation and might be leaving the domain of stable pulsation before becoming carbon-rich stars. Indeed, F. L. Schöier & H. Olofsson (2001), in a study of a large sample of bright carbon stars, have shown that only a quarter of them pulsate as Mira variables.
The naive interpretation that the preceding subsection might have suggested, namely that stars would evolve across the Mno sample, from 0.4 down to 0.2, before switching to the Myes, and later S, spectral types, is clearly flawed. Thirteen curves, of which six of spectral type S, have 0.25, larger than that of the 5 sample of Figure 2. This suggests therefore that when evolving along the TP-AGB, stars will have experienced sufficient TDU at different stages of their evolution. This is indeed what should be expected: stars of low initial masses spend their whole AGB life as Mno types and their curves may span down to 0.2 as they expand and become cooler. But stars with larger initial masses switch to the Myes spectral type earlier, before their curves reach low values of . Figure 3 shows that S types appear already at values as large as 0.27 while Myes types reach down to 0.16 and large periods without switching to S types. In this context, we note that the seven curves that have been selected for having a large period, in excess of 470 days, are of the M spectral type without known Tc content and have indeed a low value of , smaller than 0.22. The broad range of values covered by S type curves, namely by stars that are close to becoming carbon stars, is challenging interpretation. As a consequence, the selection 0.3, which we used to define the present Table 3 sample may be unjustified: some stars may leave the Mno spectral type when their curves have still a high value of , possibly as high as 0.4 in some cases. In order to progress in the understanding of this issue, it is therefore necessary to study the sample of curves having 0.3, which we do in the next subsection.
| Name | (days) | Spctr | Tc | 12/13C | (mag) | (mag) | (mag) | ||
| RT Cyg | 190 | M2-8.8 | dbfl | 1.5 | 4.7 | 0.38 | 0.54 | 0.54 | |
| T Aqr | 202 | M2-5.5 | no | 1.4 | 5 | 0.34 | 0.4 | 0.5 | |
| X Peg | 201 | M2-5 | no | 1.3 | 4 | 0.39 | 0.34 | 0.5 | |
| T Her | 165 | M2.5-8 | no | 1.5 | 5 | 0.34 | 0.44 | 0.52 | |
| T Col | 226 | M3-6 | no | 1.7 | 4.3 | 0.33 | 0.37 | 0.51 | |
| T Eri | 253 | M3-5 | no | 1.9 | 4.8 | 0.31 | 0.44 | 0.51 | |
| RY Oph | 150 | M3-6 | no | 1.7 | 4.7 | 0.32 | 0.47 | 0.52 | |
| R Boo | 224 | M3-8 | no | 1.7 | 5.1 | 0.31 | 0.35 | 0.52 | |
| R Vul | 137 | M3-7 | no | 1.8 | 4.5 | 0.37 | 0.45 | 0.51 | |
| R Vir | 146 | M3.5-8.5 | no | 1.8 | 4.3 | 0.37 | 0.34 | 0.5 | |
| U Oct | 303 | M4-6 | no | 1.8 | 5.3 | 0.27 | 0.4 | 0.53 | |
| R Oph | 303 | M4-6 | no | 2.1 | 5.6 | 0.27 | 0.43 | 0.56 | |
| R Sgr | 269 | M4-6 | no | 1.5 | 4.9 | 0.31 | 0.44 | 0.53 | |
| RR UMa | 231 | M4 | no | 1.6 | 4.8 | 0.29 | 0.44 | 0.55 | |
| R Tri | 266 | M4-8 | no | 1.7 | 5.4 | 0.25 | 0.39 | 0.52 | |
| S Lac | 240 | M4-8.2 | no | 2 | 5 | 0.32 | 0.34 | 0.53 | |
| T UMa | 256 | M4-7 | no | 2.4 | 5.3 | 0.31 | 0.62 | 0.58 | |
| R Aql | 274 | M5-9 | no | 8 | 2.3 | 4.7 | 0.31 | 0.33 | 0.53 |
| X CrB | 241 | M5-7 | no | 1.7 | 5 | 0.27 | 0.36 | 0.53 | |
| R Del | 286 | M5-6 | no | 2 | 4.8 | 0.26 | 0.38 | 0.57 | |
| R Cnc | 362 | M6-9 | no | 17 | 1.9 | 4.4 | 0.25 | 0.33 | 0.53 |
| RU Hya | 332 | M6-8.8 | no | 3.3 | 5.4 | 0.24 | 0.59 | 0.6 | |
| R Peg | 377 | M6-9 | no | 10 | 2.3 | 5 | 0.24 | 0.49 | 0.56 |
| W Peg | 345 | M6-8 | no | 17 | 2 | 4.2 | 0.28 | 0.37 | 0.52 |
| RS Vir | 354 | M6-8 | no | 3.2 | 5.5 | 0.28 | 0.53 | 0.63 | |
| R Cas | 432 | M6-10 | no | 12 | 2.1 | 6 | 0.21 | 0.84 | 0.61 |
| U Her | 404 | M6.5-9.5 | no | 19 | 2.5 | 4.9 | 0.21 | 0.4 | 0.58 |
| R LMi | 373 | M6.5-9.0 | no | 12 | 2.7 | 5.7 | 0.24 | 0.61 | 0.6 |
| R Leo | 312 | M7-9 | no | 10 | 1.2 | 4.3 | 0.26 | 0.4 | 0.56 |
| W Eri | 376 | M7-9 | no | 2.4 | 5.4 | 0.22 | 0.5 | 0.59 |
| 4 | 4 or 5 | 5 | |
| rms() | 0.350.03 | 0.290.02 | 0.240.03 |
| rms() [days] | 18937 | 26724 | 36733 |
| rms() [mag] | 4.60.3 | 5.10.3 | 5.10.6 |
| rms() [mag] | 0.410.06 | 0.410.08 | 0.510.14 |
| rms() | 0.510.01 | 0.540.02 | 0.580.03 |
| rms() [mag] | 1.60.2 | 1.90.3 | 2.40.6 |
3.3 Oxygen-rich stars having possibly experienced strong TDU for which 0.3
The curves selected in the preceding subsection were required to have 0.3. In order to see whether such a criterion is justified, we need to study the curves of the oxygen-rich stars having 0.3, excluding of course the curves of Mno type listed in Table 1, and see how different they are from those listed in Table 3. They are listed in Table 4 and Figure 4 displays their distribution in the parameter space. At first glance, Figure 4 seems to suggest that the 0.3 curves are successors of the Mnoi-j curves having 4, while the 0.3 curves are successors of the Mnoi-j curves having 3; moreover, the location of the 0.3 curves between the Mno and carbon-rich samples may suggest that the stars having 0.3 curves will evolve to the carbon-rich variables of the present sample. A close inspection of the properties of the curves listed in Table 4 is therefore necessary to understand whether such crude impressions are or not justified.
The twelve curves having 0.4 include the seven stars of spectral type M and five of the eight stars of spectral type S which the sample contains and can be interpreted as successors of the Mnoi-j curves having 4: they have =0.34 compared with 0.35, =334 compared with 189, =4.7 compared with 4.6, =0.32 compared with 0.41, =0.51 compared with 0.51 and =1.4 compared with 1.6. Namely, while the period has much increased, and a little decreased, the other parameters remained nearly unchanged on average. This is compatible with a future evolution toward the sample of carbon-rich stars as long as it would make room for a significant decrease of the oscillation amplitude.
This leaves three additional curves having 0.4, all of stars of spectral type S: R Cam, T Cam and S UMa. They are part of Sample B of Paper I. Such a small sample does not allow for drawing reliable conclusions. Yet, as can be seen on Figure 4, while the periods are compatible with an evolution toward the sample of carbon-rich stars, the values of and are surprisingly small and those of surprisingly large.
| Name | (days) | Spctr | Tc | 12/13C | (mag) | (mag) | (mag) | ||
| R And | 410 | S3.5-8.8 | yes | 40 | 2.6 | 7.4 | 0.22 | 0.76 | 0.61 |
| W And | 398 | S6-9 | yes | 1.8 | 6.3 | 0.22 | 0.79 | 0.58 | |
| S Cas | 613 | S3-5 | 32 | 3.6 | 5.6 | 0.2 | 0.58 | 0.6 | |
| Chi Cyg | 408 | S6-9 | yes | 36 | 2.1 | 8.3 | 0.18 | 0.64 | 0.56 |
| R Cyg | 427 | S2.5-6 | yes | 29 | 2.9 | 6.2 | 0.24 | 1.14 | 0.62 |
| U Cas | 277 | S3-8 | 1.2 | 6.2 | 0.25 | 0.38 | 0.53 | ||
| W Aql | 483 | S3-6 | yes | 26 | 3.2 | 5.2 | 0.26 | 0.77 | 0.62 |
| WY Cas | 479 | S6.5 | yes | 3.2 | 6.1 | 0.27 | 0.56 | 0.62 | |
| X And | 345 | S2-5 | 2.2 | 5.5 | 0.27 | 0.51 | 0.62 | ||
| SZ Cep | 329 | S3.5-4 | 1.8 | 5.7 | 0.27 | 0.36 | 0.59 | ||
| R Lyn | 378 | S2.5-6 | yes | 1.4 | 5.7 | 0.26 | 0.44 | 0.56 | |
| RR And | 328 | S6.5 | 1.3 | 5.9 | 0.24 | 0.47 | 0.57 | ||
| S Cyg | 323 | S2.5 | 0.8 | 5.2 | 0.25 | 0.5 | 0.47 | ||
| T Sgr | 391 | S4.5-5.5 | yes | 1.5 | 4.2 | 0.27 | 0.37 | 0.54 | |
| R Aur | 457 | M6.5-9.5 | yes | 33 | 2.1 | 5.9 | 0.21 | 0.5 | 0.46 |
| RU Her | 486 | M6-9 | yes | 25 | 2.3 | 5.9 | 0.18 | 0.74 | 0.54 |
| R Ser | 355 | M5-9 | yes | 14 | 2 | 6.5 | 0.24 | 0.58 | 0.57 |
| U Ari | 372 | M4-7.5 | yes | 2.1 | 6.3 | 0.23 | 0.62 | 0.59 | |
| V CMi | 366 | M4-10 | yes | 2.3 | 5.8 | 0.22 | 0.78 | 0.65 | |
| R Hor | 404 | M5-8 | yes | 2.1 | 7.3 | 0.23 | 0.53 | 0.59 | |
| S Vir | 377 | M6-9.5 | yes | 2 | 5.7 | 0.24 | 0.36 | 0.55 | |
| V Gem | 275 | M4-8 | yes | 1.5 | 5.7 | 0.26 | 0.4 | 0.56 | |
| S Ori | 421 | M6.5-9.5 | yes | 45 | 1.8 | 4.7 | 0.27 | 0.37 | 0.48 |
| V Cam | 521 | M7 | 3.6 | 5.9 | 0.17 | 0.72 | 0.61 | ||
| X Cep | 540 | M4-7 | 3.2 | 6 | 0.16 | 0.68 | 0.58 | ||
| UX Cyg | 568 | M4-6.5 | 4.4 | 4.9 | 0.2 | 0.5 | 0.62 | ||
| V Del | 534 | M4-6 | 3 | 5.2 | 0.17 | 0.88 | 0.49 | ||
| Z Pup | 512 | M4-9 | dbfl | 3.2 | 5.8 | 0.18 | 0.57 | 0.61 | |
| Z Cas | 497 | M7 | 2.7 | 4.2 | 0.19 | 0.57 | 0.66 | ||
| RU Aur | 470 | M7-9 | 3.6 | 5 | 0.21 | 0.73 | 0.59 | ||
| S CrB | 360 | M6-8 | poss | 2.8 | 5.5 | 0.26 | 0.52 | 0.64 | |
| R UMa | 301 | M3-9 | 2.9 | 5.5 | 0.27 | 0.43 | 0.61 | ||
| CH Pup | 500 | M7 | 3.4 | 4.9 | 0.23 | 0.95 | 0.6 | ||
| R CVn | 329 | M6-9 | poss | 2.1 | 4.4 | 0.3 | 0.41 | 0.51 | |
| V Cas | 229 | M5-8.5 | 1.6 | 4.8 | 0.28 | 0.28 | 0.51 | ||
| R Dra | 247 | M5-9 | dbfl | 2.1 | 5.1 | 0.3 | 0.46 | 0.55 | |
| R Tel | 462 | M5-7 | 2.5 | 5.2 | 0.16 | 0.68 | 0.53 |
The trend toward low colour indices is already apparent in the Table 4 sample of the twelve curves having 0.4. Indeed, out of the 548 curves selected by P. Merchán Benítez et al. (2023), 40 have [3.4]-[22]1.25, of which 19 are of S spectral type, 18 of M spectral type and 3 of C spectral type. Of these, 10 are part of the Table 4 sample and only 5, all of S spectral type with between 0.24 and 0.27, are part of the much larger Table 3 sample. This is evidence for a group of stars of very low colour indices, mostly of S, but also Myes spectral types, having curves of large parameter, to own a specific identity that must be associated with well-defined physical properties. A likely interpretation is offered by O. Hashimoto et al. (1998) in the case of R Hya, a star of the Table 4 sample having an M6- 9 spectral type, a colour index 0.9 and a broad curve profile, 0.36. The claim is that mass-loss has recently nearly stopped, causing the circumstellar envelope to expand away from the star and become too cold for the silicate dust grains to significantly emit in the infrared. Recent observations and analyses (L. Decin et al., 2008; A. Baudry et al., 2023; M. Joyce et al., 2024; R. Zhao-Geisler et al., 2012) have confirmed the validity of this interpretation and suggested that the cause of the absence of significant mass-loss is the likely occurrence of a thermal pulse some 200 years ago. Models such as by Y. A. Fadeyev (2024) quote a large stellar mass of 4 to 5 solar masses.
More generally, as remarked by P. Marigo et al. (2020), low mass loss rates, and therefore low colour indices, are expected when the C/O ratio is close to unity. In such a case, the amount of atmospheric oxygen when C/O1 (S type) or of atmospheric carbon when C/O1 (C type) is too low to allow for significant formation of dust on which radiation pressure could produce the mass-losing wind.
In summary, an interpretation of the Table 4 curves as associated with large initial mass stars near the end of the TP-AGB, having light curves evolving from early Mno type (4) to the carbon-rich Table 5 sample is plausible but far from being firmly established and raises several questions that deserve careful consideration.
| Name | (days) | Spctr | Tc | 12/13C | (mag) | (mag) | (mag) | ||
| R Cam | 270 | S2-8 | 0.4 | 4.5 | 0.41 | 0.28 | 0.52 | ||
| T Cam | 376 | S4-8.5 | yes | 31 | 0.7 | 5.4 | 0.4 | 0.18 | 0.49 |
| S UMa | 226 | S0-5 | yes | 1.2 | 3.7 | 0.46 | 0.09 | 0.5 | |
| V Cnc | 272 | S0-7 | yes | 1.4 | 5 | 0.33 | 0.3 | 0.55 | |
| FF Cyg | 328 | S6 | 1 | 4.5 | 0.34 | 0.22 | 0.5 | ||
| Z Del | 305 | S5.2-7 | yes | 1.4 | 5.5 | 0.3 | 0.45 | 0.51 | |
| R Gem | 361 | S2-8 | yes | 22 | 2 | 6.2 | 0.31 | 0.52 | 0.64 |
| T Gem | 287 | S1.5-9 | yes | 0.5 | 5.4 | 0.38 | 0.21 | 0.5 | |
| R Hya | 386 | M6-9 | yes | 26 | 0.9 | 3.7 | 0.36 | 0.38 | 0.49 |
| T Cep | 338 | M5-9 | yes | 33 | 1.2 | 4.1 | 0.33 | 0.29 | 0.47 |
| S Her | 307 | M4.8-7.5 | yes | 1.1 | 5.4 | 0.31 | 0.33 | 0.49 | |
| T Cas | 445 | M6-9 | prob | 33 | 1.9 | 4.3 | 0.39 | 0.27 | 0.43 |
| U UMi | 324 | M6-8 | 1.6 | 3.3 | 0.37 | 0.24 | 0.5 | ||
| S CMi | 333 | M6-8 | yes | 18 | 1.9 | 4.9 | 0.32 | 0.33 | 0.49 |
| Z Peg | 326 | M6-8.5 | yes | 1.7 | 4.7 | 0.3 | 0.33 | 0.5 |
3.4 Carbon-rich stars
The light curves of the carbon-rich stars of the present sample are listed in Table 5 and their distribution in the parameter space is illustrated in Figure 4. The width parameter exceeds 0.3. A clear correlation is seen between and and a weaker one between and and and . The analyses presented in the preceding subsections have left open a reliable identification of their progenitors. They found plausible that it hosts members of the Table 4 sample. But they could not exclude that it would also host members of the Table 3 sample. In particular, transition from Table 3 to Table 5 might occur via a regime of unstable pulsation: such curves have lost memory of the region where they come from: both =0 and =1 curves are curves of constant luminosity.
J. Bergeat & L. Chevallier (2005) have collected data related to a large sample of carbon-rich stars, of which 15 are part of the Table 5 sample. Figure 5 displays the dependence of the parameters and of their light curves on effective temperature and mass loss rate . If one excludes four low curves that are close to the domain covered by the Table 4 sample, V CrB, U Cyg, RZ Peg and R Ori, tends to increase with temperature and decrease with mass loss rate while displays the opposite trend. Such a behaviour would therefore suggest that the stars having large light curves evolve in the direction of decreasing when becoming cooler. This trend is confirmed by the dependence of on the spectral index for spectral types C: for 6, namely warmer stars, 0.530.07 while for 6, namely for cooler stars, 0.390.04. This is however excluding R For, a clear outlier having a C4 spectral type and 0.35.
To the extent that we can take as granted that carbon stars keep cooling down (P. Marigo, 2002), this suggests that the stars of the Table 5 sample having curves of large value enter it from the large side and evolve toward smaller values of . If such is the case, they would probably have a Table 3 origin: the larger , the smaller the oscillation amplitude (Figure 4), namely the closer from a regime of unstable pulsation. Figure 6 illustrates what such a transition would look like by displaying projections on the vs , vs , vs and vs planes of the distributions in parameter space of the Table 3 and Table 5 samples. It shows that, apart from and , a transition from Table 3 to Table 5 is reasonably continuous in the other parameters. The discontinuity in amplitude is not surprising: absorption of the dust envelope is known to be larger for carbon-rich rather than oxygen-rich stars.
RS Cyg and V Oph, both parts of the Figure 5 sample, have large values of , 0.66 and 0.52, respectively. They would then be likely successors of the Table 3 sample. RS Cyg has a high temperature of 3100 K and a low mass loss rate of 210-7 solar masses per year (R. Guandalini et al., 2006). Its light curve has been extensively studied and commented by R. R. Cadmus (2022, 2024). It often displays a hump on the ascending branch, close to maximum light, possibly producing a double-maximum profile. V Oph has been the target of detailed VLT observations in the infrared (K. Ohnaka et al., 2007; G. Rau et al., 2019; J. Hulberg et al., 2025) suggesting an effective temperature of 2700 K while Figure 5 quotes a temperature of 3010 K and a mass loss rate of 1.410-7 M⊙yr-1. However, the identification of the origin of curves having lower values of turns out to be difficult. BH Cru and LX Cyg were shown by S. Uttenthaler et al. (2011, 2016) to have recently switched from spectral type S to C: one would naively expect to find them near an edge of the domain covered by the carbon-rich sample in the parameter space; but such is not the case, their curves have values of (0.47 and 0.46, respectively) that sit in the middle of those of the Table 5 sample, a bit on the high side: they may be successors of the Table 3 sample as well as of the Table 4 sample, although the value of the temperature (3000 K, A. A. Zijlstra et al., 2004) rather favours the former.
G. Rau et al. (2017) and C. Paladini et al. (2017) have published a detailed comparison of the infrared spectra of R Lep and R Vol obtained by VLTI/MIDI. As the former has a value of 0.42 while the latter has the smallest value in the sample (0.31), one might expect such a comparison to reveal significant differences between the two stars. But it does not: the analyses presented by G. Rau et al. (2017) produce instead very similar stellar parameters: a luminosity of 8000-8500 solar luminosities and an effective temperature of 2900 K. R Lep is part of the Figure 5 sample with a mass loss rate of 1.610-6 M⊙yr-1. Both R Lep and R Vol are close to RZ Peg, R Ori and U Cyg, which have similar temperatures and low values.
| Name | (days) | Spctr | Tc | 12/13C | (mag) | (mag) | (mag) | ||
| AZ Aur | 415 | C8 | yes | 2.5 | 3.3 | 0.37 | 0.6 | 0.55 | |
| R Cap | 342 | C? | 3 | 3.5 | 0.35 | 0.44 | 0.57 | ||
| AX Cep | 396 | C? | 2.7 | 2.9 | 0.39 | 0.56 | 0.55 | ||
| S Cep | 484 | C7 | yes | 224 | 2.2 | 2.4 | 0.4 | 0.43 | 0.46 |
| V CrB | 358 | C6 | yes | 2.1 | 3 | 0.37 | 0.52 | 0.58 | |
| U Cyg | 463 | C9 | yes | 14 | 2.1 | 3.1 | 0.39 | 0.45 | 0.51 |
| V Cyg | 420 | C7 | 20 | 2.8 | 3.4 | 0.38 | 0.49 | 0.56 | |
| T Dra | 422 | C6 | yes | 24 | 3 | 3 | 0.37 | 0.34 | 0.56 |
| R For | 387 | C4 | 2.7 | 3 | 0.35 | 0.53 | 0.51 | ||
| T Lyn | 409 | C7 | 2 | 2.8 | 0.44 | 0.45 | 0.55 | ||
| R Ori | 377 | C8 | 1.8 | 3.5 | 0.32 | 0.27 | 0.55 | ||
| BH Cru | 513 | C8 | yes | 8 | 1.5 | 2.9 | 0.47 | 0.33 | 0.5 |
| LX Cyg | 581 | C? | yes | 32 | 1.7 | 3.4 | 0.46 | 0.41 | 0.59 |
| W Cas | 408 | C7 | yes | 25 | 1 | 2.8 | 0.45 | 0.21 | 0.52 |
| UV Aur | 395 | C8 | yes | 2.8 | 2.6 | 0.43 | 0.4 | 0.59 | |
| V Aur | 353 | C6 | 1.6 | 2.6 | 0.45 | 0.22 | 0.48 | ||
| RZ Peg | 437 | C9 | yes | 9 | 2.3 | 4.1 | 0.33 | 0.25 | 0.56 |
| RU Vir | 437 | C8 | yes | 3.7 | 3 | 0.37 | 0.72 | 0.56 | |
| R Vol | 451 | C? | 3.4 | 2.9 | 0.31 | 0.63 | 0.5 | ||
| VX Gem | 379 | C9 | yes | 9 | 1.6 | 4 | 0.36 | 0.32 | 0.59 |
| S Cam | 330 | C6 | 14 | 1.1 | 1.9 | 0.56 | 0.15 | 0.49 | |
| X Cas | 424 | C5 | 19 | 1.5 | 1.9 | 0.52 | 0.26 | 0.45 | |
| RV Cen | 445 | C3 | 1.6 | 2.4 | 0.51 | 0.24 | 0.43 | ||
| RS Cyg | 419 | C5.5 | 1.2 | 1.6 | 0.66 | 0.12 | |||
| WX Cyg | 411 | C6 | yes | 5 | 1.5 | 2.6 | 0.52 | 0.21 | 0.54 |
| U Lyr | 452 | C4 | 23 | 2.1 | 1.8 | 0.54 | 0.46 | 0.48 | |
| V Oph | 297 | C4 | yes | 11 | 1.5 | 2.4 | 0.52 | 0.29 | 0.5 |
| Y Per | 249 | C5 | 1.4 | 1.9 | 0.55 | 0.22 | 0.45 | ||
| PQ Cep | 443 | C6 | 2.2 | 2.6 | 0.46 | 0.46 | 0.47 | ||
| R Lep | 441 | C7 | 34 | 2.5 | 2.5 | 0.42 | 0.59 | 0.47 | |
| BD Vul | 430 | C6 | 1.7 | 2.3 | 0.45 | 0.27 | 0.48 | ||
| R CMi | 340 | C0 | yes | 11 | 1.3 | 2.9 | 0.47 | 0.19 | 0.57 |
S Cep, with 0.40, has an effective temperature of 2100 K (W. Nowotny et al., 2005, 2010), leaving open both a Table 3 and a Table 4 origin. Such is indeed the case for all curves having values of this order or smaller. The 13 curves having 0.4 cover rather uniformly the parameter space, failing to reveal a clear separation between two groups of curves. These examples illustrate the difficulty of producing reliable interpretations of the evolution of the stars having curves in the Table 5 sample: the likely coexistence of Table 3 and Table 4 successors is a major complication and a reliable distinction between them is difficult.
Another source of complication is the large fraction of carbon-rich stars in an unstable regime of pulsation, implying that their light curves are often excluded from our sample and cannot be parameterized reliably in the form used in the present article. P. Marigo et al. (2020) have produced a very insightful description of the evolution of carbon-rich stars along the AGB, paying attention to the initial mass of the star, the TDU efficiency, the C/O ratio and other quantities of utmost relevance. In particular, they have shown the presence of a kink in the initial-to-final mass relation, occurring at initial masses just below two solar masses. They suggest (see their Figure 6 in Supplementary information) that its progenitors spend a fraction of their carbon star life as semi-regular variables, characterised by small amplitude pulsations, low C/O ratio, small mass-loss rates and low terminal velocities. A detailed study, case by case, of the curves of Table 5 should be made within the framework of their study; however, this is well beyond the scope of the present article.
Finally, we recall that we have shown in Paper I the dependence of and on and remarked that carbon-rich stars and oxygen-rich stars follow similar trends, the former having slightly lower values of and of than the latter for a same value of . Figure 7 illustrates this issue in the present sample, displaying separately the samples of Tables 1, 3, 4 and 5. The Table 1 distributions illustrate the transition from what was defined as first and second families of Mno curves in Paper I, with both and increasing; the location of the curve of R Cas (0.84) confirms the suggestion that was made in Subsection 3.2 that the star is in the process of switching to the Myes spectral type. The same trend is observed in the distributions of the Table 3 sample, with a global shift toward larger and ; a clear shift toward large values of gives evidence for the increased instability of the pulsation regime. The concentration in small values of both and illustrates the difference between the curves of the Table 4 sample and the others. The distributions displayed by the Table 5 sample confirm the comments made in Paper I.
3.5 Humpy curves and hump-less curves
In Paper I, we remarked that when a hump was observed on the ascending branch, there was a relation between its location and the parameters , and : namely the closer it was from minimal light, the larger were the values of these three parameters. Figure 8 illustrates the present situation. We selected curves displaying a clear hump on the ascending branch and classified them as type b, c or d according to the hump location: b for humps close to minimal light, d for humps close to maximal light and c in between. In the definitions of both what we call a strong hump and of which class the curve belongs to, there is of course some subjectivity which must be kept in mind. Moreover, what we define as humps are rather abrupt changes of the rate of increase of the luminosity: while c types are real humps, b and d types are rather flattening of the light minimum and maximum, respectively. Figure 8 reveals a number of remarkable features: all d types are from the Table 4 sample and, with the exception of V Cyg, all b types are from the Table 3 sample; instead, c types are present in all three samples and span a very large range of values; the general trends observed in Paper I are clearly confirmed and complemented with the dependence of the hump type on , and . Such humps may simply be the trivial effect of non-linear dynamics, similar humps are observed in RR Lyrae stars and in bump Cepheids with the Hertzsprung progression. In such a case, they may not be amenable to a simple interpretation, such as tracking a shock wave as suggested by L. S. Kudashkina & G. M. Rudnitskij (1994). Yet, their distribution in the parameter space suggests that they carry some additional information, which we however fail to decrypt.
4 Other stars
4.1 Hot Bottom Burners
R Cen and R Nor have been identified as clear Hot Bottom Burning (HBB) candidates at the beginning of the TP-AGB by D. A. Garc´ıa-Hernández et al. (2013) and S. Uttenthaler et al. (2011, 2012) on the basis of detailed evaluations of the abundances of technetium, lithium, rubidium and zirconium in the spectra: they are Tc-poor and Li-rich and have low mass-loss rates at the level of 10-7 M⊙ yr-1, properties which they share with two other clear HBB candidates, SV Cas and RU Cyg, both semi-regular variables. They are expected to have initial masses of 4 to 6 solar masses, in contrast to more massive HBB candidates with very high mass-loss rates (10-5 M⊙ yr-1), known as OH/IR stars, at the end of the TP-AGB (D. A. Garc´ıa-Hernández et al., 2006, 2007, 2009). Their light curves are double-peaked (Figure 9) and that of R Cen has recently experienced a sudden change in period and amplitude. Moreover, they are spectacular outliers in colour vs period diagrams (S. Uttenthaler et al., 2019; P. Merchán Benítez et al., 2023; P. Whitelock et al., 2000). They have been commented upon in numerous earlier studies, both for R Cen (P. C. Keenan & J. Landi Dessy, 1966; G. Hawkins et al., 2001; W. S. G. Walker & J. Greaves, 2001; M. R. Templeton et al., 2005; J. R. Garcia & S. A. Dominguez, 2006; N. Vogt et al., 2016; B. Rea & J. C. Martin, 2024) and for R Nor (N. Vogt et al., 2016). Most authors take it as granted that the sudden changes observed in the light curve of R Cen are the result of a recent thermal pulse, as had been first proposed by P. R. Wood & D. M. Zarro (1981). Measurements of the radial velocity in the near infrared (T. Lebzelter et al., 2005) have given evidence for a clear anti-correlation with the light curve.
The light curve of SV Cas, rather than being double-peaked, displays a hump on the descending branch, which is exceptional. The amplitude of the oscillation increased from 1.5 mag at the end of the 1980s to 3 mag today. The mass loss rate has been measured as 2.410-7 M⊙ yr-1 (P. Scicluna et al., 2022) and the 12C/13C ratio as 102 (K. H. Hinkle et al., 2016). Over the past century, the light curve of RU Cyg (R. R. Cadmus, 2021, 2024) has evolved from a relatively regular sequence of double-peaked cycles to more irregular cycle profiles; while the period has stayed at its value of 468 days, the amplitude of the oscillations has decreased from nearly 2 to 0.5 units of magnitude. Both SV Cas and RU Cyg light curves (Figure 9) share with those of R Cen and R Nor the relative brevity of the ascending and descending branches, implying a large value of . We noted that another star displays a light curve that resembles those of the four clearly identified HBB candidates: DH Cyg (Figure 9). The star has been studied by G. H. Smith & E. A. Bashay (2014) as a metal-poor field giant and its emission in the mid-infrared has been found to be very strong, together with AC Her. The recent changes experienced by the R Cen curve and the changing regime of pulsation displayed by the curves of SV Cas and RU Cyg are evidence for HBB stars at the beginning of the TP-AGB to be close to the limits of the domain of stable pulsation.
The double-peaked profiles of R Cen and R Nor are outstanding. The and parameters do not obey the standard relation between them, the former being not as large as implied by the latter because of the depression between the two maxima. We looked for possible other cases among Long Period Variables but did not find any convincing case. An example is shown in the second panel of Figure 9 that compares mean profiles of curves sharing with R Cen and R Nor the property of having both a large value of the width parameter, , and a small value of the colour index, . The curves of BH Cru, LX Cyg and S Cep, carbon-rich stars of which the light curves are occasionally quoted as double-peaked, as are those of TT Cen, RS Cyg and U CMi, are of a different nature: they simply display a hump on the ascending branch. So does the curve of RU Tau, a Mira variable of spectral type M3.5 quoted by P. Merchán Benítez et al. (2023) as having a very large meandering period of 589 days and a colour index 1.9; it is the only oxygen-rich Mira variable sharing with R Cen and R Nor a large period, low colour indices and a large width (0.41), and as such deserves being given careful attention; however, the profile of its light curve favours a Table 4 rather than HBB assignment.
In contrast with the light curves of AGB stars, those of RV Tauri stars commonly display double-peaked cycles of a shape similar to those displayed by R Cen and R Nor (Figure 9). RV Tauri stars (R. Manick et al., 2019; I. Gezer et al., 2015, and references therein) are thought to be metal-deficient, population II, low mass post-AGB stars, often surrounded by a dust disc, generally associated with the presence of a Main Sequence companion, which causes periodic obscuration of the light curve. Their periods range between 30 and 150 days and, having nearly lost their circumstellar envelope, they have low masses, typically between 0.3 and 0.9 solar masses. Numerous detailed studies of the shapes of the light curves (Y. Tuchman et al., 1993; A. B. Fokin, 1994; J. R. Buchler et al., 1995; Z. Kollath et al., 1998) have shown that they display a broad range of irregularity. Whether the understanding of the dynamics governing the pulsations of an archetypal RV Tau star, AC Her (Figure 9), may help or not with the understanding of the dynamics governing the pulsations of R Cen and R Nor is unclear. On the one hand, the similarity between the light curves is so striking that one feels that the dynamics must share some common features. On the other hand, one may fear that such similarity is the simple result of the basic properties of non-linear dynamics, which can be expressed in apparently simple terms using the language of fundamental and overtone modes, but which fail to tell much about the actual physical mechanisms at stake (A. B. Fokin, 1994), which, in the cases of AC Her vs R Cen/R Nor, are obviously very different.
4.2 Mno outliers
Five stars of Mno spectral type were excluded from the sample listed in Table 1 as having light curves presenting peculiarities that deserve a special discussion, W Dra, W Hya, T Ari, U CMi and T UMi.
The curve of W Dra is only marginally different from those of the Table 1 sample, in which it could be included: we did not find in the published literature sufficient reason to give it a special treatment.
The curve of W Hya is instead very different from those of the Table 1 sample and closely resembles those of some stars of spectral type Myes listed in Table 4, such as R Hya. S. Uttenthaler et al. (2011) give a detailed description of the search for technetium in the spectra of R Hya and W Hya, both displaying significant period changes. They conclude positively for the former and negatively for the latter but note that intrinsically weak lines may go entirely undetected if a Mira star is observed at a pulsation phase with severe line weakening. Being nearby, W Hya has been the target of a very large number of high-resolution observations, each revealing detailed features that could not have been detected if the star were more distant. Recent examples quoting relevant references are D. T. Hoai et al. (2022b) and T. Khouri et al. (2020). The curve of T Ari is similar to curves of typical carbon-rich stars such as WX Cyg. Together with that of U CMi, its profile has a very large parameter, 0.5, which is otherwise found only for carbon-rich stars or Hot Bottom Burners. Unfortunately, too little is known about these stars to suggest plausible explanations and the cases must remain unexplained.
Finally, T UMi is well-known as a case of sudden period change, with a decline from more than 300 days to about 200 days over the past 20 years, causing it to leave the domain of stable pulsation. Before this change, the parameters of the light curve were as listed in Table 6, indeed very different from those of Table 1. J. A. Mattei & G. Foster (1995), J. Gal & K. Szatmary (1995) and K. Szatmáry et al. (2003) suggested that T UMi is currently just before the beginning of a helium shell flash; S. Uttenthaler et al. (2011) remarked that, in such a case, the lack of Tc and Li in the spectrum imply that T UMi is either not massive enough to experience a TDU event, or that the thermal pulse is early and/or weak. Indeed, both Y. A. Fadeyev (2022) and L. Molnár et al. (2019) have recently produced state-of-the-art models consistent with such an interpretation, showing that T UMi had an initial mass between 1.0 and 1.5 solar masses (1.20.2 for Molnár et al.). This makes T UMi very different from the Hot Bottom Burning R Cen and illustrates how similarity of light curves cannot be naively interpreted as similarity of underlying dynamics.
4.3 Binaries
Cet and R Aqr are well-known Mira variables of spectral type Myes, which were not retained in the selection of P. Merchán Benítez et al. (2023) as being members of binaries. Their curve parameters are listed in Table 6 and, apart from a slightly too low oscillation amplitude for R Aqr, they fit well in the group of curves listed in Table 3, suggesting that binarity does not have a major impact on the arguments developed in the present article.
X Oph is a Mira variable of spectral type Mno3-9 that had not been retained in Paper I for having a light curve too different from the other Mno stars studied in the article. Indeed, its curve parameters listed in Table 6 display much too low an oscillation amplitude in comparison with the curves listed in Table 1. A likely explanation is the presence of a bright K-type giant companion (J. D. Fernie, 1959), which, being unresolved, would be the source of the light detected in the minima of the light curve.
4.4 Other stars
In the present subsection we briefly comment about the light curves of other stars, which we came across on the occasion of the present study and which may deserve some mention.
- R Car (A. Rosales-Guzmán et al., 2023, 2024) was recently the target of high resolution observations. The parameters describing its light curves are listed in Table 6, compatible with both the Table 1 and Table 3 samples. Both in terms of luminosity, temperature and size (M. Takeuti et al., 2013) and in terms of light curve (Table 6), R Car is very similar to Cet suggesting that it has a modest initial mass and is marginally dredging-up material from the helium burning shell.
- U Per is a Mira variable that was used by M. S. Cunha et al. (2020), together with L2 Pup, to illustrate general considerations that they make on the stability of the pulsation regime of variable stars. The light curve of U Per (Table 6) belongs clearly to the Table 4 sample, close to the domain covered by carbon-rich stars.
- V Hya is a carbon star thought to be close to transiting to post-AGB (L. Planquart et al., 2024b, and references therein), with a high mass-loss rate of 10-5 M⊙yr-1 (G. R. Knapp et al., 1999) and a mean spectral broadening of 13.5 km s-1with a periodic variation of 9 km s-1 suggesting rapid star rotation (C. Barnbaum et al., 1995). Numerous observations suggest the presence of an unseen companion at 11 au distance causing periodic dimming (17 years) of the light curve (L. Planquart et al., 2024a, and references therein). R. Sahai et al. (2022) have used ALMA to detect the emission of CO lines with a good angular resolution and R. Zhao-Geisler et al. (2012) have used VLTI/MIDI to observe the dust and molecular shells surrounding the star and compare their properties with those of four AGB stars studied in the present article: R Aql, R Aqr, R Hya and W Hya. Both the periodic dimming and an insufficient coverage offered by existing observations prevent a reliable evaluation of the parameters of the light curve, which is anyhow far from a regime of stable pulsation. Yet, we selected cycles allowing for a crude estimate of their values, which are listed in Table 6. They fit well in the large region of the Table 5 sample.
- VX Sgr is a very massive and luminous oxygen-rich AGB star which has been the target of a multitude of observations (A. Chiavassa et al., 2009; H. M. Tabernero et al., 2021, and references therein), which are still challenging interpretation. Its light curve is outside the domain of stable pulsation and its complex nature prevents a meaningful evaluation of the curve parameters defined in the present article. Yet, it does experience episodes of relatively stable pulsation, a particularly long one occurring between 1968 and 1997. During this episode, it is possible to estimate approximate values of the curve parameters, which are listed in Table 6. If it were not for the period, which exceeds 2 years, larger than any of our sample, the other parameters would clearly assign the light curve to Table 4, near the border with the Table 5 curves of carbon-rich stars and significantly shifted in their direction.
| Name | (days) | Spctr | Tc | 12/13C | (mag) | (mag) | (mag) | ||
| HBB and HBB-like | |||||||||
| R Cen | 531 | M4-6 | no | 2.2 | 4.2 | 0.34/0.63 | 0.23 | ||
| R Nor | 498 | M3-7 | no | 2.3 | 4.8 | 0.28/0.53 | 0.35 | ||
| SV Cas | 450 | M5 | no | 2.7 | 2.2 | 0.29/0.42 | |||
| RU Cyg | 468 | M9 | no | 32 | 2.4 | 1.7 | 0.43/0.78 | ||
| DH Cyg | 531 | M6 | 2.7 | 2.8 | 0.36/0.72 | ||||
| AC Her | 75.4 | F4 | 7 | 1.1 | 0.55/0.87 | 0.08 | |||
| Mno outliers | |||||||||
| W Dra | 279 | M3-4 | no | 2.7 | 4.8 | 0.35 | 0.5 | 0.59 | |
| W Hya | 391 | M7.5-9 | no | 16 | 0.4 | 3.2 | 0.37 | 0.34 | 0.48 |
| T Ari | 320 | M6-8 | no | 15 | 1.6 | 2.1 | 0.5 | 0.22 | 0.55 |
| U CMi | 409 | M4 | no | 2 | 4 | 0.47 | 0.34 | 0.39 | |
| T UMi | 317 | M6 | no | 1.4 | 4.9 | 0.21 | 0.31 | 0.53 | |
| Binaries | |||||||||
| R Aqr | 386 | M6.5-8.5 | yes | 16 | 1.6 | 4.2 | 0.23 | 0.6 | 0.57 |
| Cet | 333 | M5-9 | yes | 10 | 1.5 | 5.6 | 0.24 | 0.62 | 0.62 |
| X Oph | 334 | M3-9 | no | 12 | 2 | 1.7 | 0.35 | 0.22 | 0.49 |
| Other cases | |||||||||
| RU Tau | 589 | M3.5 | 1.9 | 3.7 | 0.41 | ||||
| U Ori | 372 | M6-9.5 | poss | 25 | 2.6 | 5.5 | 0.24 | 0.55 | 0.61 |
| R Car | 308 | M5-8 | 1.4 | 5 | 0.25 | 0.47 | 0.53 | ||
| U Per | 321 | M5.5-7 | prob | 1.6 | 3.2 | 0.47 | 0.22 | 0.46 | |
| V Hya | 530 | C6-7-N | 2.5 | 2 | 0.62 | 0.25 | 0.5 | ||
| VX Sgr | 759 | M8 | 2 | 4 | 0.46 | 0.56 | 0.51 | ||
5 Discussion and conclusion
The emphasis in the present article has been to try to reveal general relations between the evolution of stars along the AGB and the parameters of their light curves, at the price of accepting some loss of rigour. For example, by using mean parameters , and , we lost the information contained in the differences between the members of the relevant pairs. Also, we accepted to ignore the differences, often significant, between the uncertainties attached to the parameters that have been used. This must be kept in mind when considering the results obtained and possibly extending the present work to future observations or analyses. By using five parameters to characterize the main features of a light curve, we have been able to clarify the picture that had been sketched in Paper I. Figure 10 and Table 7 illustrate the main results, which we summarize below. While on the early part of the AGB, stars expand, cool down and tend to pulsate with increasing periods; when their period exceeds 100 days, they are classified as Mira variables and enter the sample considered in the present work. At that time, they are oxygen-rich with an M spectral type and have not yet experienced sufficient TDU events for their impact to be detectable. The sample of Mno curves listed in Table 1 corresponds to such stars. As the temperature of the star decreases and as its size increases, the light curve evolves from what was defined as first Mno family to what was defined as second Mno family in Paper I. Namely (black arrow in Figure 10) the period increases, the width parameter decreases, the phase at minimal light increases, weak humps may appear on the ascending branch close to minimal light, the pulsations become less regular, the colour indices increase and the ratio between the oscillation amplitudes in the infrared and in the visible increases.
| (days) | (mag) | (mag) | (mag) | |||
| Table 1 | 27479 | 0.290.05 | 4.90.5 | 2.00.5 | 0.440.11 | 0.540.04 |
| Table 3 | 41092 | 0.230.04 | 5.80.8 | 2.40.8 | 0.560.20 | 0.570.04 |
| Table 4 | 32652 | 0.350.05 | 4.70.8 | 1.30.5 | 0.290.10 | 0.510.04 |
| Table 5 | 41061 | 0.440.08 | 2.80.6 | 2.10.7 | 0.390.15 | 0.530.04 |
What the present work has made clear is that the transition from Mno to Myes, namely the occurrence of TDU events strong enough to have a detectable impact, occurs at very different stages along the TP-AGB. Indeed it is expected to never occur for stars of low initial masses (F. Herwig, 2005), and the curves of such stars stay in the Table 1 sample for the whole AGB life of the star. The presence in the Table 3 sample of stars of Myes, and even S spectral types covering a broad spectrum of values is compatible with the expectation that TDU events occur the earlier on the TP-AGB, the larger the initial mass of the star. The curves of the Table 3 sample (blue arrow in Figure 10) seem to keep evolving in the same direction after the transition to the Myes and S spectral types: the periods and amplitudes of the oscillations keep increasing and they become less regular. This is consistent with the expectation that thermal pulses modify only progressively the internal state of the star. The absence of carbon-rich curves in the parameter space of the curves close to the more evolved curves of the Table 3 sample suggests that some of these curves leave the domain of stable pulsation and become chaotic before the transition from oxygen-rich to carbon-rich.
Another major result obtained by the present work is the existence of curves leaving the Table 1 sample very early and transiting to the Myes and S spectral types with the width parameter increasing rather than decreasing. These are the curves of the Table 4 sample and their location in the parameter space suggests that they will evolve to the Table 5 sample when the stars will become carbon-rich (red arrows in Figure 10). Here again, such a suggestion, albeit sensible and plausible, is far from being proven. The dispersion of the curve parameters about their mean is large for both the Table 4 and Table 5 samples, showing the complexity of the situation. In particular, the presence in the Table 5 sample of curves evolving with decreasing (possible successors of the Table 3 sample) and of curves evolving with increasing (possible successors of the Table 4 sample) contribute to such complexity. It is natural to speculate that the stars of the Table 4 sample have larger initial masses, and probably are closer to exhausting their convective layers than the stars of the Table 3 sample: the behaviour of the curves of HBB stars discussed in Subsection 4.1 gives support to such speculation as do the curves of stars that are known to be close to post-AGB, such as V Hya and VX Sgr.
A major difficulty that has been recurrently met during the study is the lack of a clear picture of the physics governing the transition from a stable to an unstable regime of pulsation, namely of the nature of the stability strip in the Hertzsprung-Russell diagram. The distinction between Mira, semi-regular and irregular variables, like the distinction between sudden, continuous and meandering period changes, are tools that prove useful in organizing the discussion, but they fail to provide a clear picture of the physics at stake. An abundant literature is dedicated to what causes changes of pulsation regime, most of the time blaming the occurrence of a thermal pulse, but little is known about how such changes actually develop. They clearly have to be seen in the context of non-linear dynamics, with light curves prone to display features such as humps or double-peaking and to become chaotic, but we are unable to answer simple questions such as: will R Dor, a nearby AGB star that has been extensively observed and studied (T. R. Bedding et al., 1998), see its light curve evolve from its current irregular regime to a more stable one? and if yes, when will that happen?
We spotted a few cases that raise questions: DH Cyg seems to be a good HBB candidate and its spectrum deserves being inspected accordingly. The light curve of W Hya suggests that the star is Tc-rich rather than Tc-poor, possibly revealing that it already experienced significant TDU events, however not strong enough to produce a detectable technetium signal. T Ari is known to have a Tc-poor spectrum but its light curve displays the features usually displayed by the curves of carbon-rich stars; together with U CMi, its light curve profile has a very large parameter, challenging interpretation. T UMi is another well-known puzzling case of a Tc-poor star and the inspection of its light curve confirms its peculiarity. Finally, inspection of its light curve has also confirmed the well- known oddity of the behaviour of VX Sgr.
In summary, the present study has significantly clarified the picture that had been sketched in Paper I. In particular it has offered a detailed description of the relation between stellar parameters and the properties of the light curves of the Table 1 (Mno) sample and it has revealed a clear distinction between the curves of the Table 3 and Table 4 samples, which had been previously unnoticed. It has suggested that their evolution follows two different paths in the parameter space, the former possibly implying a transition out of the regime of stable pulsation. However, it has raised many questions which remain unanswered and the interpretations that it has suggested are far from being ascertained. It should serve as a guideline for future observations and analyses, which would hopefully help with progressing with the general understanding of the issues that have been addressed. It makes one feel that light curves still carry messages and information that we have not yet been able to decrypt.
acknowledgments
To the extent that the analyses presented in the present article may have contributed some progress, the credit belongs to the innumerable observers around the world and to the AAVSO, without whom none of the results could have been obtained. We are accordingly deeply indebted to them: to the many observers for the high quality of their observations and to the AAVSO for the outstanding handling and reduction of the data that makes their use particularly easy and efficient. We are deeply grateful to Pr Tim Bedding for his interest in our work and for useful comments. Financial support from the World Laboratory, the Rencontres du Vietnam and the Vietnam National Space Center is gratefully acknowledged. This research is funded by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 103.99-2024.36.